If a segment with a slope of two passes through points (2, 1) and (z, 9), what is the value of z?
(A) -14 (B) -2 (C) 3 (D) 6 (E) 8

Answer by richwmiller(17219)

(y-y)/(x-x)=m
(9-1)/(z-2)=2
8=2*(z-2)
4=z-2
6=z
(D)=6